A glass container was initially charged with 1.70 moles of a gas sample at 3.75 atm and 21.7 °C. Some of the gas was released as the temperature was increased to 28.1 °C, so the final pressure in the container was reduced to 0.998 atm. How many moles of the gas sample are present at the end?
Answer #1

Given:
Pi = 3.75 atm
Pf = 0.998 atm
ni = 1.70 mol
Ti = 21.7 oC
= (21.7+273) K
= 294.7 K
Tf = 28.1 oC
= (28.1+273) K
= 301.1 K

use:
(Pi*Vi)/(ni*Ti) = (Pf*Vf)/(nf*Tf)
(3.75 atm)/(1.7 mol*294.7 K) = (0.998 atm)/(nf*301.1 K)
nf = 0.4428 mol
Answer: 0.443 mol